UT 20171102 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt (LSW), DThompson (LBTO)
Observer AO-Team: DMiller, GTaylor (LBTO)
Telescope Operator: SAllanson

Observations done remotely from Tucson

Plan for the night

- Time is shared with LBTB

- LUCI1/2 AFC AO-ima commissioning in binocular mode for half a night

- If seeing is bad during first part of night, PEPSI-tests will be executed for about one hour (LBTB reimbursed later on)

- Otherwise if seeing is bad LBTB get up to half a night, more bad time goes to LBTO

Summary

Details

All times are UT
Files are luci1/2.20171102.NNNN.fits

00:34 Sunset, Weather: clear skies, T = 3deg, hum = 40%, wind ~ 5-10 m/s

00:50 Steve checks pointing and collimates,

Doug: AO-man @ SX
Greg: AO-man @ DX

Seeing looks good, we will start

01:00 Preset to YSO2153, preset OK, too early (too bright)...

01:10 Preset to YSO2153 again, preset OK, RunAO, no RS found, Steve checks pointing

01:18 Preset again, again no RS, target coords seem to be wrong

DIMM reports ~ 1"

AFC at medium elevation

01:22 Preset to M39, preset OK, RunAO, failed on SX (human error), loop closed, bin = 2, 200Hz (SX), 400 Hz (DX)

at the same time issues with the compensation mirror service - can't turn on

passive flexure compensation (tab goes always back to "no flexure corrections")

Roll back to LUCI-SW version we used on Oct 28, didn't help.

Called Alexander to fix it and he fixed it.
Problem: The service hung because of an aborted SRR-process
This is a major bug and needs to be removed.

01:55 start with AFC-testing...finally...

DIMM ~ 0.8"

Did a set in Ks and Br_gam at ele 65-74, OK

L1: 11-35, L2: 12-36

AFC at high elevation

04:07 Preset to Mayall 2, Steve checks pointing before, preset OK, RunAO,
SX OK, DX error just after closing (camera lense moved) , corrected, OK,

DIMM ~ 0.9

Ks: L1: 41-43 (SRR), 44-48 (short), L2: 41-43 (SRR), 44-48 (short), ele: 80-83

DIMM ~ 1.1, seeing starts to degrade

05:07 Started an image to align AFC-holes for H-band ima.

Aborted SRR-images on LUCI1 and LUCI2, compensation mirror service on LUCI2 hangs,
same problem as earlier tonight, stop/restart SW, worked

No DIMM as we are too high in ele

05:22 cont. with H-band

stopped after 1st ima since AFC@L1 went bananas, restarted,

H: L1: 59-61(SRR), 62-66 (short), L2: L2: 61-63 (SRR), 64-68 (short), ele=81-78

skip frame at images 65/67 due probably due to seeing spike

J: L1: 75-77 (SRR), L2: 75-77 (SRR), 78-82 ele: 77-71

stopped after 1st ima since AFC@L2 went bananas,

Seeing back to 0.9" but fluctuating, 2 loop openings due to seeing spikes, skipped short images
last set of images taken with Fe II

FeII: L1: 80-83 (SRR), 84-88 (short), L2: 80-83 (SRR), 84-88 (short), ele 71-65, OK

loop opened at SX again, not to seeing , ima 82 crap

End of half night

Did the first successfull tests with AFC-spec on sky afterwards using the 1.6687mue OH-line in
parallel to the LBTB -observer using MODS smile
Topic revision: r5 - 02 Nov 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback